OpenStudy (anonymous):

How many different three letter permutations can be formed from the letters in the word clipboard?

OpenStudy (callisto):

Number of letters = 26 Since, they are different => no repetition Permutation required = 26 x 25 x 24 =?

OpenStudy (anonymous):

i got 729, what do u think

OpenStudy (callisto):

Disagree...

OpenStudy (kinggeorge):

I'm getting 504. You have 9 letters in the word "clipboard" and they are all distinct. Thus, the number of 3 letter words is \[9\cdot8\cdot7=504\]

OpenStudy (kropot72):

The correct answer is 9 * 8 * 7 = 504 from the formula:\[nPr=\frac{n!}{(n-r)!}\]

OpenStudy (callisto):

Sorry.. I read the question wrong... my bad!!! :(

OpenStudy (anonymous):

does the permutations mean multiple times which it is 9^3 = 729 or there are 9 letters and they are all different, so if you can use each letter only once in each permutation* it is 9 x 8 x 7 = 504. this is where I get stuck

OpenStudy (kinggeorge):

You can only use each letter once, which is why it's 9x8x7=504 instead of 9^3=729

OpenStudy (anonymous):

ok thanks everybody ; )

OpenStudy (kinggeorge):

You're welcome.